50 grams of ice-cold water (0 °C) is heated to boiling (100 °C). The specific heat capacity of water is 4.18 J /g°C. How much energy did the water absorb?


1
Expert's answer
2021-11-13T01:33:11-0500

"Q= m.c.\u2206T"


Q=heat absorbed

"m=mass=50grams"

"c" = specific heat capacity "=418J\/g\u00b0C"

"\u2206T=100-0=100\u00b0C"



"Q=50\u00d74.18\u00d7(100-0)=20900J"